# Non-commutative Iwasawa theory of abelian varieties over global function fields

### v1

@@ 7. The order of vanishing of the characteristic element
Since Proposition 7.6 tells us that the characteristic series of $H^{i}(H_{\rho},A_{p^{\infty}}(F_{\infty}))$ does not contribute a factor of $T$, we see from (7.3) that the same conclusion holds for $H^{i}(H_{\rho},\mathrm{tw}_{\rho}(A_{p^{\infty}})(F_{\infty}))$. ∎
Given any infinite separable extension $K$ of $F$ contained in $F_{S}$, we can define the twisted Selmer group ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/K)$ to be
${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/K)=\ker\left(H^{1}(G_{S}(K),\mathrm{tw}_{\rho}(A_{p^{\infty}}))\longrightarrow\bigoplus_{v\in S}J_{v}(\mathrm{tw}_{\rho}(A)/K)\right).$
Here, for each $v\in S$,
@@
where $L$ runs over all finite extensions of $F$ contained in $K$, $w$ runs over all primes of $L$ lying above $v$ and the limit is taken with respect to the restriction maps. We shall write $X(\mathrm{tw}_{\rho}(A)/K)$ for the Pontryagin dual of ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/K)$.
@@
We have the following short exact sequence
$0\longrightarrow{\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F_{\infty})\longrightarrow H^{1}(G_{S}(F_{\infty}),\mathrm{tw}_{\rho}(A_{p^{\infty}}))\longrightarrow\bigoplus_{v\in S}J_{v}(\mathrm{tw}_{\rho}(A)/F_{\infty})\longrightarrow 0$
of cofinitely generated $\Lambda_{\mathcal{O}}(H)$-modules.
@@
By the definition of $\rho$, we see that $G_{S}(F_{\infty})$ acts trivially on $W_{\rho}$, and so we have ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F_{\infty})={\mathrm{Sel}}(A/F_{\infty})\otimes_{\mathcal{O}}W_{\rho}$. Since ${\mathrm{Sel}}(A/F_{\infty})$ is cofinitely generated over $\Lambda(H)$ by Lemma 3.3, it follows that ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F_{\infty})$ is cofinitely generated over $\Lambda_{\mathcal{O}}(H)$. Via similar arguments, we see that $H^{1}(G_{S}(F_{\infty}),\mathrm{tw}_{\rho}(A_{p^{\infty}}))$ and $\bigoplus_{v\in S}J_{v}(\mathrm{tw}_{\rho}(A)/F_{\infty})$ are cofinitely generated over $\Lambda_{\mathcal{O}}(H)$. Since $W_{\rho}$ is free over $\mathcal{O}$, the short exact sequence of the lemma can be obtained from Theorem 4.2 by tensoring the short exact sequence in the said theorem with $W_{\rho}$. ∎
By a similar argument to that in [10, Proposition 2.5], the restriction map
@@
has kernel which is cofinitely generated over $\mathcal{O}$. It then follows from this and the preceding lemma that ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F^{\mathrm{cyc}})$ is cofinitely generated over $\mathcal{O}$. Therefore, we may apply the structure theory for $\Lambda_{\mathcal{O}}(\Gamma)$-modules to obtain the following pseudo-isomorphism
@@
where the $f_{j}$’s are irreducible distinguished polynomials.
We can now state the semi-simplicity conjecture of Greenberg for Artin twists of abelian varieties over a global function field.
@@
Let $A$ be an abelian variety defined over a global function field $F$. Then the action of $\Gamma$ on $X(\mathrm{tw}_{\rho}(A)/F^{\mathrm{cyc}})$ is completely reducible, that is, $\beta_{j}=1$ for all $j$ in (7.4).
@@
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Let $\rho$ be an irreducible Artin representation of $G={\mathrm{Gal}}(F_{\infty}/F)$ and let $\xi_{A}$ denote a characteristic element of $X(A/F_{\infty})$. Suppose that Hypothesis $(\mathbf{G_{\rho}})$ holds. Then we have
@@
In addition, equality holds when Greenberg’s semi-simplicity conjecture 7.11 holds for $\mathrm{tw}_{\rho}(A)/F$.
@@
The arguments used to prove Theorem 7.7 carry over for the Artin twists, taking Proposition 7.9 and Lemma 7.10 into account, and noting that $X(\mathrm{tw}_{\rho}(A)/F_{\infty})=\mathrm{tw}_{\hat{\rho}}X(A/F_{\infty})$ (cf. [7, Lemma 3.4]). ∎
@@ 8. The Generalised Euler characteristics
In this section, we give the definition of the generalised Euler characteristic. The idea of defining a finite Euler characteristic $\chi(G,Y)$ for a discrete $p$-primary $G$-module $Y$ when the cohomology groups $H^{i}(G,Y)$ are infinite was first conceived by Coates–Schneider–Sujatha in the form of the truncated Euler characteristic [9]. Our presentation here will follow that of Zerbes [40].
@@
For a discrete $p$-primary $G$-module $Y$, we define
@@
where the middle map is $\phi_{Y^{H}}$ and the last map is given by inflation. Similarly, for $i\geq 1$, define
@@
where the first map is given by restriction and the middle map is equal to $\phi_{H^{i}(H,Y)}$. Note that the first map is surjective since ${\rm cd}_{p}(\Gamma)=1$. We define $d^{-1}$ to be the zero map.
All the above properties for the morphisms involving in the definition of the complex $(H^{i}(G,Y),d^{i})$ will become clear in the proof of the following proposition.
@@
We have the Hochschild–Serre spectral sequence is given by
@@
Since ${\rm cd}_{p}(\Gamma)=1$, we have $H^{r}(\Gamma,H^{s}(H,Y))=0$ unless $r=0$ or $1$. Thus, there exists a short exact sequence
(8.1) $0\to H^{1}(\Gamma,H^{s-1}(H,Y))\to H^{s}(G,Y)\to H^{0}(\Gamma,H^{s}(H,Y))\to 0$
for all $s\geq 1$. Now, we see that $d^{s}:H^{s}(G,Y)\to H^{s+1}(G,Y)$ is equal to the composition
@@
where the surjective map and injective map are exactly given by the short exact sequence (8.1). Therefore, $d^{s+1}\circ d^{s}$ must factor through the following composition
@@
which is the exact sequence (8.1), and so must be the zero map. Hence, we have $d^{s+1}\circ d^{s}=0$, as desired. ∎
@@
Denote by ${\mathfrak{H}}^{\bullet}$ the cohomology of the complex in Proposition 8.1. If ${\mathfrak{H}}^{i}$ is finite for all $i\geq 0$, we say that the $G$-module $Y$ has finite generalised $G$-Euler characteristic $\chi(G,Y)$, defined by
@@
    If the groups $H^{i}(G,Y)$ are finite for all $i$, the generalised $G$-Euler characteristic for $Y$ agrees with the usual $G$-Euler characteristic for $Y$.
@@
    If $H^{0}(G,Y)$ and $H^{1}(G,Y)$ are infinite and $H^{i}(G,Y)$ is finite for all $i\geq 2$. The module $Y$ has finite generalised $G$-Euler characteristic if and only if the map $d^{0}$ has finite kernel and cokernel, in which case
@@
This is essentially established in the work of Zerbes [40] but not explicitly stated in the form here. For the convenience of the readers, we shall supply a proof of the lemma. For the first assertion, it suffices to show that for a cochain complex $({\mathfrak{X}}^{\bullet},\delta^{\bullet})$ with ${\mathfrak{X}}^{i}$ finite for all $i\geq 0$, we have
@@
To see this, consider the following two short exact sequences
@@
$0\to B^{i}\to Z^{i}\to H^{i}({\mathfrak{X}}^{\cdot})\to 0$
As all modules in the sequences are finite, we have
#$\#({\mathfrak{X}}^{i})=\#(Z^{i})\cdot\#(B^{i+1}).$
Therefore, noting that #$\#(B^{0})=1$, we have
$\prod_{i\geq 0}\#({\mathfrak{X}}^{i})^{(-1)^{i}}=\prod_{i\geq 0}(\#(Z^{i})\cdot\#(B^{i+1}))^{(-1)^{i}}=\prod_{i\geq 0}\#(H^{i}({\mathfrak{X}}^{\cdot}))^{(-1)^{i}},$
and (1) follows. To prove the second assertion, we first note that $\ker(d^{0})={\mathfrak{H}}^{0}$. Thus, if $d^{0}$ has finite kernel and cokernel, then ${\mathfrak{H}}^{0}$ is finite. Now, from the exact sequence
@@
where the third map is induced by $d^{1}$, we see that ${\mathfrak{H}}^{1}$ is also finite. As ${\mathfrak{H}}^{i}$ are finite for $i\geq 2$ by assumption, $Y$ has finite generalised $G$-Euler characteristic. Conversely, suppose that $Y$ has finite generalised $G$-Euler characteristic. Since $\ker(d^{0})={\mathfrak{H}}^{0}$, we see that $d^{0}$ has finite kernel. Since $H^{2}(G,Y)$ is finite, $\ker(d^{2})$ is finite. Now, considering the exact sequence
$0\to{\mathfrak{H}}^{1}\to\mathrm{coker}~{}\!d^{0}\to\ker d^{2}\to{\mathfrak{H}}^{2}\to 0,$
we see that $\mathrm{coker}~{}\!(d^{0})$ is finite. It remains to show the formula (8.2). The complex $(H^{\bullet}(G,Y),d^{\bullet})$ induces the following cochain complex
@@
Since all terms in the cochain complex are finite, noting that $\ker(\bar{d}^{1})\simeq{\mathfrak{H}}_{1}$, we see from (1) that
@@
Then noting that #$\#({\mathfrak{H}}^{0})=\#(\ker(d^{0}))$ and dividing both sides by the equality (8.3), we arrive at the following equality
@@
which is (8.2), as required. ∎
We now record the following important result which relates the generalised Euler characteristic to Akashi series.
@@
Let $Y$ be a discrete $p$-primary $G$-module such that $Y^{\vee}\in{\mathfrak{M}}_{H}(G)$. If $Y$ has finite generalized $G$-Euler characteristic, then the leading term of its Akashi series $\mathrm{Ak}(Y)$ is $\alpha_{Y}T^{k}$, where $k$ is equal to the alternating sum
@@
See [40, Proposition 2.10]. ∎
We note that it is in general not true that the generalised $G$-Euler characteristic is multiplicative in short exact sequences (see [40, Remark after Lemma 2.13]). However, the following lemma suffices for most of our purposes.
@@
Let $0\to B_{1}\to B_{2}\to B_{3}\to B_{4}\to 0$ be an exact sequence of discrete $G$-modules with the property that $B_{1}$ and $B_{4}$ have finite $G$-Euler characteristics. Then $B_{2}$ has finite generalised $G$-Euler characteristic if and only if $B_{3}$ does.
@@
Set $C=\ker(B_{3}\to B_{4})$. By separately considering the following two short exact sequences
$0\to B_{1}\to B_{2}\to C\to 0$
@@
we may assume that either $B_{1}=0$ or $B_{4}=0$. We shall prove the case when $B_{4}=0$, the case of $B_{1}=0$ being similar. Thus, we may suppose that we have a short exact sequence of $G$-modules
$0\to B_{1}\to B_{2}\to B_{3}\to 0$
with $B_{1}$ having finite $G$-Euler characteristics. In particular, we have that $H^{i}(G,B_{1})$ is finite for every $i\geq 0$. Taking the long exact sequence of $G$-cohomology of the short exact sequence above, we see that
@@
have finite kernel and cokernel for all $i\geq 0$. By the naturalness of the Hochschild–Serre spectral sequence and our definition of $d^{i}$, we see that $\psi_{i}$ are cochain maps between the complexes $(H^{\bullet}(G,B_{2}),d^{\bullet}_{B_{2}})$ and $(H^{\bullet}(G,B_{3}),d^{\bullet}_{B_{3}})$. Since the kernels and cokernels of these cochain map are finite, a direct homological algebraic argument shows that the induced maps on their cohomology also have finite kernel and cokernel. Hence it follows that $B_{2}$ has finite generalised Euler characteristic if and only if $B_{3}$ does. ∎

### v2

@@ 7. The order of vanishing of the characteristic element
Since Proposition 7.6 shows that the characteristic series of $H^{i}(H_{\rho},A_{p^{\infty}}(F_{\infty}))$ does not contribute a factor of $T$, it follows from (7.3) that the same holds for $H^{i}(H_{\rho},\mathrm{tw}_{\rho}(A_{p^{\infty}})(F_{\infty}))$. ∎
Given any infinite separable extension $K$ of $F$ contained in $F_{S}$, we define the twisted Selmer group ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/K)$ by
${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/K)=\ker\left(H^{1}(G_{S}(K),\mathrm{tw}_{\rho}(A_{p^{\infty}}))\longrightarrow\bigoplus_{v\in S}J_{v}(\mathrm{tw}_{\rho}(A)/K)\right),$
where for each $v\in S$,
@@
Here, the direct limit is taken over all finite extensions $L$ of $F$ contained in $K$, and $w$ runs over all primes of $L$ lying above $v$. The transition maps in the limit are given by restriction. We denote by $X(\mathrm{tw}_{\rho}(A)/K)$ the Pontryagin dual of ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/K)$.
@@
We have the following short exact sequence of cofinitely generated $\Lambda_{\mathcal{O}}(H)$-modules
$0\longrightarrow{\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F_{\infty})\longrightarrow H^{1}(G_{S}(F_{\infty}),\mathrm{tw}_{\rho}(A_{p^{\infty}}))\longrightarrow\bigoplus_{v\in S}J_{v}(\mathrm{tw}_{\rho}(A)/F_{\infty})\longrightarrow 0.$
@@
By the definition of $\rho$, the group $G_{S}(F_{\infty})$ acts trivially on $W_{\rho}$. Hence, we have ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F_{\infty})={\mathrm{Sel}}(A/F_{\infty})\otimes_{\mathcal{O}}W_{\rho}$. Since ${\mathrm{Sel}}(A/F_{\infty})$ is cofinitely generated over $\Lambda(H)$ by Lemma 3.3, it follows that ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F_{\infty})$ is cofinitely generated over $\Lambda_{\mathcal{O}}(H)$. Similarly, we know that
$H^{1}(G_{S}(F_{\infty}),\mathrm{tw}_{\rho}(A_{p^{\infty}}))\quad\textrm{and}\quad\bigoplus_{v\in S}J_{v}(\mathrm{tw}_{\rho}(A)/F_{\infty})$
are cofinitely generated over $\Lambda_{\mathcal{O}}(H)$. Since $W_{\rho}$ is a free $\mathcal{O}$-module, the short exact sequence of the lemma then follows from Theorem 4.2 by tensoring the corresponding short exact sequence in that theorem with  $W_{\rho}$. ∎
By an argument similar to that in [11, Proposition 2.5], the restriction map
@@
has kernel which is cofinitely generated over $\mathcal{O}$. It then follows from this fact and the preceding lemma that ${\mathrm{Sel}}(\mathrm{tw}_{\rho}(A)/F^{\mathrm{cyc}})$ is cofinitely generated over $\mathcal{O}$. Therefore, we can apply the structure theory for $\Lambda_{\mathcal{O}}(\Gamma)$-modules to obtain the following pseudo-isomorphism
@@
where each $f_{j}$ is an irreducible distinguished polynomial.
The following is Greenberg’s semi-simplicity conjecture for Artin twists over global function fields, which serves as the function field analogue to the conjecture in [27, Conjecture 6.4].
@@
Let $A$ be an abelian variety defined over a global function field $F$ with ${\rm char}(F)\neq p$. Then the action of $\Gamma$ on $X(\mathrm{tw}_{\rho}(A)/F^{\mathrm{cyc}})$ is completely reducible, that is, $\beta_{j}=1$ for all $j$ in (7.4).
@@
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Let $\rho$ be an irreducible Artin representation of $G={\mathrm{Gal}}(F_{\infty}/F)$, and let $\xi_{A}$ denote a characteristic element of $X(A/F_{\infty})$. Suppose that condition $(\mathbf{G_{\rho}})$ holds. Then
@@
Moreover, equality holds if Greenberg’s semi-simplicity conjecture 7.11 holds for $\mathrm{tw}_{\rho}(A)/F$.
@@
The arguments used in the proof of Theorem 7.7 carry over to the setting of the Artin twists, using Proposition 7.9 and Lemma 7.10. In the proof, we also make use of the identity $X(\mathrm{tw}_{\rho}(A)/F_{\infty})=\mathrm{tw}_{\hat{\rho}}X(A/F_{\infty})$, as shown in [8, Lemma 3.4]. ∎
@@ 8. The Generalised Euler characteristics
In this section, we introduce the definition of the generalised Euler characteristic. The idea of defining a finite Euler characteristic $\chi(G,Y)$ for a discrete $p$-primary $G$-module $Y$, even when the cohomology groups $H^{i}(G,Y)$ are infinite, was first proposed by Coates–Schneider–Sujatha in the form of the truncated Euler characteristic [10]. Our presentation here follows that of Zerbes [50].
@@
Now, let $Y$ be a discrete $p$-primary $G$-module, we define
@@
where the middle map is $\phi_{Y^{H}}$, and the last map is the inflation map. Similarly, for each $i\geq 1$, define
@@
where the first map is the restriction map, the middle map is $\phi_{H^{i}(H,Y)}$. Note that the restriction map is surjective since ${\rm cd}_{p}(\Gamma)=1$. We define $d^{-1}$ to be the zero map.
All the properties for the morphisms involved in the definition of the complex $(H^{i}(G,Y),d^{i})$ will become clear in the proof of the following proposition.
@@
We consider the Hochschild–Serre spectral sequence
@@
Since ${\rm cd}_{p}(\Gamma)=1$, it follows that $H^{r}(\Gamma,H^{s}(H,Y))=0$ unless $r=0$ or $1$. Thus, for all $s\geq 1$, we obtain the short exact sequence
(8.1) $0\to H^{1}(\Gamma,H^{s-1}(H,Y))\to H^{s}(G,Y)\to H^{0}(\Gamma,H^{s}(H,Y))\to 0.$
Now, observe that the differential $d^{s}:H^{s}(G,Y)\to H^{s+1}(G,Y)$ is given by the composition
@@
where the surjection and injection come directly from the short exact sequence (8.1). Therefore, $d^{s+1}\circ d^{s}$ factors through
@@
which, by exactness of (8.1), is the zero map. Hence, we conclude that $d^{s+1}\circ d^{s}=0$, as required. ∎
@@
Let $\mathfrak{H}$$^{\bullet}$ denote the cohomology of the complex constructed in Proposition 8.1. If ${\mathfrak{H}}^{i}$ is finite for all $i\geq 0$, we say that the $G$-module $Y$ has a finite generalised $G$-Euler characteristic $\chi(G,Y)$, defined by
@@
    If the groups $H^{i}(G,Y)$ are finite for all $i$, then the generalised $G$-Euler characteristic of $Y$ agrees with the usual $G$-Euler characteristic.
@@
    Suppose $H^{0}(G,Y)$ and $H^{1}(G,Y)$ are infinite, and $H^{i}(G,Y)$ is finite for all $i\geq 2$. Then $Y$ has a finite generalised $G$-Euler characteristic if and only if the map $d^{0}$ has a finite kernel and cokernel, in which case
@@
This result is essentially contained in the work of Zerbes [50], though not explicitly stated in this form. For the reader’s convenience, we provide a full proof. For part (1), it suffices to show that for a cochain complex $({\mathfrak{X}}^{\bullet},\delta^{\bullet})$ with each ${\mathfrak{X}}^{i}$ finite for $i\geq 0$, we have
@@
To verify this, consider the following two short exact sequences
@@
$0\to B^{i}\to Z^{i}\to H^{i}({\mathfrak{X}}^{\cdot})\to 0.$
Since all modules involved are finite, we have
$({\mathfrak{X}}^{i})=\#(Z^{i})\cdot\#(B^{i+1}).$
Because $(B^{0})=1$, it follows that
$\prod_{i\geq 0}\#({\mathfrak{X}}^{i})^{(-1)^{i}}=\prod_{i\geq 0}(\#(Z^{i})\cdot\#(B^{i+1}))^{(-1)^{i}}=\prod_{i\geq 0}\#(H^{i}({\mathfrak{X}}^{\cdot}))^{(-1)^{i}}.$
This proves part (1). For part (2), first note that $\ker(d^{0})={\mathfrak{H}}^{0}$. Thus, if $d^{0}$ has a finite kernel and cokernel, then ${\mathfrak{H}}^{0}$ is finite. Consider the exact sequence
@@
where the third map is induced by $d^{1}$. Since $\mathrm{coker}~\!d^{0}$ and $H^{2}(G,Y)$ are finite, it follows that ${\mathfrak{H}}^{1}$ is also finite. Given that ${\mathfrak{H}}^{i}$ is finite for all $i\geq 2$ by assumption, thus $Y$ has a finite generalised $G$-Euler characteristic. Conversely, suppose that $Y$ has a finite generalised $G$-Euler characteristic. Since $\ker(d^{0})={\mathfrak{H}}^{0}$ and ${\mathfrak{H}}^{0}$ is finite, $d^{0}$ must have finite kernel. Because $H^{2}(G,Y)$ is finite, so is $\ker(d^{2})$. Now consider the exact sequence
$0\to{\mathfrak{H}}^{1}\to\mathrm{coker}~\!d^{0}\to\ker d^{2}\to{\mathfrak{H}}^{2}\to 0.$
It follows that $\mathrm{coker}~\!(d^{0})$ is finite as well. To establish the formula (8.2), note that the complex $(H^{\bullet}(G,Y),d^{\bullet})$ induces the following cochain complex
@@
All terms in this complex are finite, and $\ker(\bar{d}^{1})\simeq{\mathfrak{H}}_{1}$. By part (1), we obtain
@@
Since $({\mathfrak{H}}^{0})=\#(\ker(d^{0}))$, dividing both sides of (8.3) by $({\mathfrak{H}}^{0})$ gives
@@
which is exactly the expression in (8.2). ∎
Let $|\cdot|_{p}$ denote the $p$-adic valuation of $\overline{{\mathbb{Q}}}$, normalised so that $|p|_{p}=p^{-1}$. We now record the following important result, which relates the generalised Euler characteristic to the Akashi series.
@@
Let $Y$ be a discrete $p$-primary $G$-module such that $Y^{\vee}\in{\mathfrak{M}}_{H}(G)$. If $Y$ has a finite generalised $G$-Euler characteristic, then the leading term of its Akashi series $\mathrm{Ak}(Y)$ is given by $\alpha_{Y}T^{k}$, where $k$ is the alternating sum
@@
See [50, Proposition 2.10]. ∎
We remark that, in general, the generalised $G$-Euler characteristic is not multiplicative in short exact sequences (see [50, the remark following Lemma 2.13]). However, the following lemma suffices for most of our purposes.
@@
Let
$0\to B_{1}\to B_{2}\to B_{3}\to B_{4}\to 0$
be an exact sequence of discrete $G$-modules such that $B_{1}$ is finite and $B_{4}$ has a finite $G$-Euler characteristic. Then $B_{2}$ has a finite generalised $G$-Euler characteristic if and only if $B_{3}$ does.
@@
Set $C=\ker(B_{3}\to B_{4})$. By analysing the following two short exact sequences separately:
$0\to B_{1}\to B_{2}\to C\to 0,$
and
@@
we may reduce to the case where either $B_{1}=0$ or $B_{4}=0$. We prove the lemma in the case $B_{1}=0$; the case of $B_{4}=0$ is analogous. Assume, therefore, that we have a short exact sequence of $G$-modules
$0\to B_{2}\to B_{3}\to B_{4}\to 0,$
with $B_{4}$ having finite $G$-Euler characteristics. In particular, $H^{i}(G,B_{4})$ is finite for all $i\geq 0$. Consider the long exact sequence in $G$-cohomology arising from the short exact sequence above. It yields morphisms
@@
with finite kernel and cokernel for all $i\geq 0$. By the naturality of the Hochschild–Serre spectral sequence and the definition of $d^{i}$, the maps $\psi_{i}$ are cochain maps between the complexes $(H^{\bullet}(G,B_{2}),d^{\bullet}_{B_{2}})$ and $(H^{\bullet}(G,B_{3}),d^{\bullet}_{B_{3}})$. Since the kernel and cokernel of these cochain maps are finite, standard homological algebra argument implies that the induced maps on the cohomology of these complexes also have a finite kernel and cokernel. Therefore, $B_{2}$ has a finite generalised Euler characteristic if and only if $B_{3}$ does. ∎